The yield to maturity on a bond is currently 8.20 percent. The real rate of return is 3.59 percent. What is the rate of inflation? A. 4.63 percent B. 3.53 percent C. 4.45 percent D. 3.92 percent
Answer #1

Based on Fischer Relation,

(1 + Nominal rate) = (1 + Real Rate) * (1 + Inflation)

(1 + 8.20%) = (1 + 3.59%) * (1 + Inflation)

(1 + Inflation) = 1.082/1.0359

1 + Inflation = 1.0445

Inflation = 0.0445 = 4.45%